Subsitiution method
5x + 4 =33
x=4y-3

Answer by nyc_function(2741)

5(4y - 3) + 4 = 33
20y - 15 + 4 = 33
20y - 11 = 33
20y = 33 + 11
20y = 44
y = 44/20
y = 11/5
==================
x = 4y - 3
x = 4(11/5) - 3
x = 44/5 - 3
x = 29/5
Final answer: x = 29/5 and y = 11/5